QUESTION 5 When 50mL of 0.10 M HOAc is titrated with 25mL of 0.10 M NaOH, the pH is оооо
Answer #1

Ka for acetic acid = 1.8 x 10-5 Number of moles of CH3COOH = M*V = 0.10 M* 50 mL = 5 mmol Number of moles of NaOH = M*V = 0.1